OpenStudy (anonymous):

the width, w, of a rectangular garden is x - 2. The area of the garden is x^3 - 2x -4. What is an expression for the length of the garden? a. x^2 - 2x - 2 b. x^2 + 2x - 2 c. x^2 - 2x + 2 d. x^2 + 2x + 2

OpenStudy (espex):

Have you ever done polynomial division?

OpenStudy (anonymous):

no.

OpenStudy (espex):

You know that the Area (A) of a rectangle is the Length (L) times the Width (W) or \(A=L\times W\)

OpenStudy (espex):

That means, the Length (L) is equal to the Area (A) divided by Width (W) or \(L = \frac{A}{W}\)

OpenStudy (anonymous):

okay.

OpenStudy (espex):

The problem statement gives you your width and area, can you write out the problem using the equation \(L=\frac{A}{W}\)?

OpenStudy (anonymous):

L = x - 2/x^3 -2x -4

OpenStudy (espex):

Close, you just need to flip the two, you have W/A and you want A/W.

OpenStudy (anonymous):

L = x^3 - 2x - 4/ x - 2

OpenStudy (espex):

Okay, now let's put it in a form we understand.

OpenStudy (espex):



OpenStudy (anonymous):

ok

OpenStudy (espex):

Now since we have only two items out front, let's look at what it would take to get the first two terms under the division. So how many times does x-2 go into x^3-2x

OpenStudy (anonymous):

3

OpenStudy (espex):

Well, 3 * x-2 = 3x-6, and you cannot take that away from x^3-2x.

OpenStudy (espex):

But what if you multiplied it by 'x'?

OpenStudy (anonymous):

how do i do that

OpenStudy (espex):

What do you get if you multiply x * (x-2)?

OpenStudy (anonymous):

im not sure how to do that

OpenStudy (espex):

You would multiply x*x and x*-2 to get you x^2-2x. But you cannot take away x^2 from x^3, so you will need to multiply by x^2 and try again.

OpenStudy (espex):

What would you multiply your (x-2) by to get an x^3?

OpenStudy (espex):

Okay, I'll do the first step. You have (x-2) and you are trying to find as close, without going over, as you can get to x^3-2x (because we are just looking at the first two terms.

OpenStudy (espex):